A 7 strand aluminium conductor is 1.2 miles long. Its resistance at 35 °C is 2Ω. Determine the diameter of each strand in inches.


1
Expert's answer
2021-03-30T19:54:23-0400
"R=\rho\frac{l}{nA}\\2=(2.7\cdot10^{-8})\frac{1931}{7A}\\A=3.7\cdot10^{-6}\ m^2=3.7\ mm^2"